OpenStudy (kingkas):

A movie theater sold 1000 tickets and made $7160 in one night. If an adult ticket costs $9 and a children's ticket costs $5, how many adult tickets were sold?

OpenStudy (anonymous):

...are there any answer choices? ._.

OpenStudy (kingkas):

A. a + c = 7160 9a + 5c = 1000 B. a c = 1000 9a 5c = 7160 C. a + c = 14 9a + 5c = 8160 D. a + c = 1000 9a + 5c = 7160 @aaran24

OpenStudy (kingkas):

D